# Consistency of EFT illuminated via relative entropy: A case study in scalar field theory

### v1

@@ 3.1.2 U(1) symmetry-breaking potential
    In the right panel of Fig. 2, we display the normalized relative entropy of the $j$-th order of $\lambda$ compared to that of the second order of $\lambda$ in percentage, i.e., $100\times|s^{(j)}-s^{(2)}|/|s^{(2)}|$, as a function of $\dot{\phi}_{0}/m_{2}r_{\rm min}$. The grey line indicates that its size is 100 %. The curves represent the third, fourth, fifth, and sixth orders of $\lambda$ of the normalized relative entropy compared to their second order of $\lambda$, with red, blue, green, and purple, respectively. The black dotted curve displays its numerically evaluated magnitude at an infinite order of $\lambda$. In regions where $\dot{\phi}_{0}/m_{2}r_{\rm min}$ is considerably smaller than $\sim 0.1$, each order effect is less than 100 % compared to $s^{(2)}$, indicating the validity of the perturbation. For instance, the sign of the relative entropy up to the third and fifth orders of $\lambda$ changes when the red and green curves intersect the grey line, respectively. In the regions with $\dot{\phi}_{0}/m_{2}r_{\rm min}$ larger than $\sim 0.1$, the effect of each finite order is more than 100 % compared to $s^{(2)}$, and the perturbation is invalid. Based on the analysis of these panels, it is evident that the points where the sign flips of the relative entropy occur represent the boundaries at which the operator expansion is either valid or invalid and quantify the validity of the EFT description. This fact can be understood analytically through the order-of-magnitude evaluation for $j>2$, $|S^{(j)}/S^{(j-1)}|\sim|S^{(3)}/S^{(2)}|$. According to eq. (91), $S^{(2)}$ is non-negative, and the non-negativity breaking of $s^{(3)}$ (representing $S^{(2)}+S^{(3)}\sim 0$) indicates $|S^{(j)}/S^{(j-1)}|\sim 1$ for $j>2$, as confirmed by numerical calculation. At this juncture, it becomes unfeasible to quantitatively truncate the operator expansion beyond the third order, causing a breakdown in the operator expansion (or the EFT description).
@@
  • $\tilde{\lambda}_{4}=1$ and $\dot{\phi}_{0}/m_{2}r_{\rm min}=10^{-1}$ — Figure 3 depicts a plot similar to Fig. 2, assuming $\tilde{\lambda}_{4}=1$ and $\dot{\phi}_{0}/m_{2}r_{\rm min}=10^{-1}$ with a horizontal axis of $\tilde{\lambda}_{3}-1$. Compared to Fig. 2, the relative entropy shows almost the same qualitative behavior, except for the difference in the horizontal axis. In particular, it can be seen that, as shown in Fig. 2, the regions where the perturbation is valid are characterized by the point where a sign flip in the relative entropy occurs.
    Figure 3: Left panel: Normalized relative entropies, $s^{(2)}$ (black), $s^{(3)}$ (red), $s^{(4)}$ (blue), $s^{(5)}$ (green), $s^{(6)}$ (purple), and $s^{(\infty)}$ (black dotted) as a function of $\tilde{\lambda}_{3}-1$. Right panel: The size of normalized relative entropies, $s^{(3)}$ (red), $s^{(4)}$ (blue), $s^{(5)}$ (green), $s^{(6)}$ (purple), and $s^{(\infty)}$ (black dotted) compared with $s^{(2)}$ as a function of $\tilde{\lambda}_{3}-1$, respectively. The parameters are assumed to be $\tilde{\lambda}_{4}=1$ and $\dot{\phi}_{0}/m_{2}r_{\rm min}=10^{-1}$.
@@ 3.2 Three scalar fields theory
We consider a scenario where a new heavy field generates the U(1) symmetry-breaking effect in eq. (85). Consider the following three scalar fields theory,
@@
where $\phi_{1}$ and $\phi_{2}$ are the same degrees of freedom as the scalar fields appearing in eq. (72), and $\phi_{3}$ is a real scalar field with a mass heavier than $\phi_{1}$ and $\phi_{2}$, and it behaves as a new particle in the viewpoint of the theory (72), and $V^{(3)}(\phi_{i})$ is a potential of this theory. To preserve the flatness of the inflaton potential as in eq. (73), we assume U(1) symmetry under a phase shift of $\phi_{1}+i\phi_{2}=re^{i\theta}$ in the potential $V^{(3)}(\phi_{i})$. By using the fields $r,\theta$ and $\chi:=\phi_{3}$, the action (99) can be rewritten at renormalizable level as follows:
@@
Upon substituting eq. (101) into eq. (45), the Euclidean effective action, up to the fourth order of the perturbation, is calculated as follows:
@@
where $m^{2}_{\sigma}:=m^{2}_{2}-\beta^{2}r_{\rm min}^{4}/m_{3}^{2}$ denotes the squared mass of $\sigma$. By comparing eq. (103) with eq. (86), we find that the U(1) symmetry-breaking parameters $\tilde{\lambda}_{3}$ and $\tilde{\lambda}_{4}$ are generated from the new particle $\chi$ as follows:
@@ 3.3 An EFT of inflation generated from more general UV theories
To consider the phenomenological implications of the above inequalities (110) and (111), let us imagine the following situation. Consider the case where the Wilson coefficients $c_{2}$, $c_{3}$, etc., are determined by evaluating the observables and comparing them with experimental data. If those values of the Wilson coefficients do not satisfy the above inequalities (110) and (111), considerations based on the relative entropy suggest the following possibilities: (i) the physical phenomena are not described by the EFT (65), including the breakdown of the operator expansion; (ii) the physical phenomena are described by the EFT (65), but the loop-level contribution to the Wilson coefficients dominates; or (iii) the EFT (65) describe the phenomena, but the EFT (65) is not generated through the interaction of heavy and light fields. For each of the three cases above, we mention some possibilities. In case (i), for instance, the phenomenon is described by a theory with entirely different degrees of freedom and symmetries than the EFT (65), or the operator expansion breaks down, as seen in the sections above, meaning that UV theory, which also includes heavy degrees of freedom, is needed to describe the phenomenon. In case (ii), the UV theory of EFT (65) is the strongly coupled theory, meaning that the perturbation does not represent physics. In particular, this case implies that the UV theory is the strongly coupled theory, in the same sense as the $S$-matrix-based perturbative unitarity argument. Case (iii) includes, for instance, the possibility that there is no UV theory corresponding to the EFT (65), which could lead to pathological behavior, such as unitarity violation. The fact that physical models can belong to these three cases is expected to provide valuable insights into identifying models describing physical phenomena. In the next section, we will see that there are parameter regions allowed in non-Gaussianity experimental data that suggest the possibility of (i), (ii), and (iii).
@@ 4 Non-Gaussianity and non-negativity of relative entropy
Figure 4: The observationally 68% (dark blue), 95% (blue), and 99.7% (light blue) confidence regions in the $(c_{s},\tilde{c}_{3}(c_{s}^{-2}-1))$ parameter space Planck:2019kim . In the red-shaded regions, the non-negativity of relative entropy is violated in perturbation theory, and the validity of the EFT (65) is lost, i.e., the operator or loop expansion becomes invalid when the EFT is generated by the interaction between heavy and light fields.
@@
where $\dot{\theta}_{0}:=\dot{\phi}_{0}/r_{\rm min}$, and these quantities are expressed in terms of parameters defined by the mass eigenstates of the heavy field in the theory with the interactions between the heavy and light fields. $m$ denotes the mass of the heavy particle (148) in that theory involving the interactions, and $\lambda_{3}$ is defined in eq. (149) of Appendix C. These parameters were introduced for convenient comparisons with previous studies Kim:2021pbr . For a detailed derivation of eq. (116), see Appendix C or Ref. Kim:2021pbr . Also, the power spectrum is expressed as,
@@
where, from the second condition of the weakly coupled theory defined in Sec. 2.2.1 (validity of operator expansion), we used $(c_{2}+\frac{3}{2}c_{3}\dot{\phi}_{0}^{2})^{-1}\simeq c_{2}^{-1}(1-\frac{3}{2}c_{3}\dot{\phi}_{0}^{2}/c_{2}+\cdots)$ and focused on the main contribution of the operator expansion. Figure. 4 shows allowed regions on $(c_{s},\tilde{c}_{3}(c_{s}^{-2}-1))$ plane from the current experiment Planck:2019kim . In red shaded regions in Fig. 4, the inequality (119) does not hold, and the validity of the EFT is broken. Also, by combining eqs. (115), (118) and (119), an upper bound on $f_{\rm NL}$ is obtained as follows:
@@
We see that when the EFT is generated through the weakly coupled theory (see Sec. 2.2.1 for its definition) with the interactions between heavy and light fields, the observationally allowed parameter region is limited. It should be emphasized that the red shaded regions in Fig. 4 are parameter regions that can be realized by the strongly coupled theory and is not physically inadmissible.
Figure 5: Contour plot of $f_{\rm NL}$ in the two scalar fields theory (85), shown in the plane of $m/H$ and $(\lambda_{3}-1)(\dot{\theta}_{0}/m)^{2}$, for $\dot{\theta}_{0}/H=10^{-1}$. Please also see Appendix C and Ref. Kim:2021pbr for the definition of $\lambda_{3}$ and $\lambda_{4}$. The colored dotted curves represent the perturbative unitarity bounds (123) (blue), and (124) (green). We choose $\Lambda_{\rm 5pt}=\Lambda_{\rm 6pt}=10^{5}H$ for the blue and green curves. In the red-shaded regions, the non-negativity of relative entropy is perturbatively violated, and the theory (85) behaves as the strongly coupled theory.
Before closing this section, we compare our method of characterizing the perturbation validity with conventional scattering amplitude-based perturbative unitarity results. Here, we compare the two methods, in particular, taking as an example the theory of two scalar fields in Sec. 3.1.2. First, from the inequality (97) up to the third order of the perturbation in the relative entropy and relations presented in Appendix C, we obtain
@@
This result represents that the inequality (119), which holds for the EFT of inflation generated from general UV theories, can be confirmed using a concrete example. Next, we consider inequalities arising from the unitarity of the $S$-matrix, which has been studied in previous literature Kim:2021pbr . In theories for which the perturbative calculations are valid, inequalities arising from the unitarity of the $S$-matrix should also be satisfied at the tree-level for high-energy scatterings. As studied in Ref. Kim:2021pbr , deviations of $\lambda_{3}$ and $\lambda_{4}$ [^2] Note that $\lambda_{3}$ and $\lambda_{4}$ are the parameters defined in Appendix C for the mass eigenstate of the heavy field in the theory with the interactions between heavy and light fields and are different from $\tilde{\lambda}_{3}$ and $\tilde{\lambda}_{4}$. See Appendix C for the relation between $\tilde{\lambda}_{3,4}$ and ${\lambda}_{3,4}$. from one break the perturbative unitarity of scattering amplitudes, yielding cutoff scales of the theory if the perturbation represents physics faithfully. In the theory (85), the cutoff scales calculated from 5 and 6-point scatterings are provided as Kim:2021pbr
@@
From eqs. (123) and (124), we also see that the perturbative calculation of the scattering amplitude is valid for $\lambda_{3}=\lambda_{4}=1$, even in high energy regions. This fact reminds us that the non-negativity of relative entropy is unbreakable for $\lambda_{3}=\lambda_{4}=1$. It should be noted here that the scattering amplitude method characterizes the condition under which loop-level contributions do not exceed tree-level ones, whereas the relative entropy method can, in addition, characterize the validity of the operator expansion (see Sec. 2.2 for details). The $S$-matrix and the relative entropy are quantities related to the unitarity, but due to these differences, different constraints would represent the validity of the perturbation. From eqs. (123) and (124), we obtain
@@
As these relations depend on the cutoff scale of the theory, we show these bounds in Fig. 5 for some benchmark values of the cutoff scale. Please look at the caption of Fig. 5 for more information. As shown in Fig. 5, considerations based on the relative entropy can reveal regions of the non-linear parameter where the perturbative calculations are valid and are complementary to constraints from the conventional perturbative unitarity for a given cutoff scale. It should be emphasized that while constraints from the perturbative unitarity are highly dependent on the unknown cut-off scale, the relative entropy method is not.
5 Discussion
This section discusses the assumptions used in obtaining the results mentioned above, i.e., conditions of the weakly coupled theory. As explained before, the first condition of the weakly coupled theory defined in Sec. 2 is closely related to the perturbative unitarity in the scattering amplitude. This is because the dominance of the tree-level contribution compared to the loop-level one in each operator of the effective action means that the tree-level effect is dominant in the scattering amplitude. In this work, we did not consider the loop-level contributions in evaluating the relative entropy because the emphasis is on its non-negativity in the weakly coupled theory. For instance, if the UV theory includes heavy fermions in addition to scalar fields, assuming the weakly coupled theory means that loop-level contributions arising from the fermion fields are negligible. However, our results of this study might not be applicable when the main contribution in the EFT is at the loop-level, such that the heavy degrees of freedom in the UV theory consist only of fermion fields. As studied in Ref. Cao:2022ajt ; Cao:2022iqh , there are classes of UV theories to which the results of this study can be adapted at the loop-level, while there are also classes of UV theories to which they cannot be adapted [^3] More specifically, there are theories for which the third term in the last line of eq. (7) is not zero at the loop-level. Such theories are not considered in this study, and the results of this study may not be available.. The calculations of the relative entropy at the loop-level have been studied in Ref. Cao:2022ajt ; Cao:2022iqh ; please look at the details there. Assuming the weakly coupled theory also implies the validity of saddle-point approximation in the path integral, as mentioned in Ref. Cao:2022ajt ; Cao:2022iqh . This is because classical-level evaluations are inadequate when performing Euclidean path integrals near maxima points of the action, and higher-order loop-level corrections are essential.
Let us now discuss the second condition of the weakly coupled theory. We set the condition for the validity of the operator expansion to be able to truncate higher-order operators at arbitrary orders quantitatively. For instance, the perturbation in the following case is valid, although, at first glance, it does not meet this validity condition. Suppose that, up to a particular order $n$-th of the operator, the higher-order contributions of the operators are more significant than the lower-order contributions, but when the order is bigger than $n$-th, the contributions of the higher-order operators are sufficiently smaller than those of the lower-order operators. In this case, the higher-order operator cannot be quantitatively truncated at any order smaller than $n$-th, but the perturbation can be regarded as working well for sufficiently high-order perturbative calculations. However, the perturbations in this case would not be valid because it is essential to evaluate them to all orders, as the specific order $n$ can only be determined by considering the contributions from all perturbative orders. In the present work, the weakly coupled theory is defined so that situations requiring infinite-order evaluation of the perturbation are not included in the valid perturbation regions.

### v5

@@ 3.1.1 U(1) symmetric potential
    In the right panel of Fig. 2, we display the normalized relative entropy of the $j$-th order of $\lambda$ compared to that of the second order of $\lambda$ in percentage, i.e., $100\times|s^{(j)}-s^{(2)}|/|s^{(2)}|$, as a function of $\dot{\phi}_{0}/m_{2}r_{\rm min}$. Here, the $s^{(2)}=1$ from eq. (97). The grey line indicates that its size is 100 %. The curves represent the third, fourth, fifth, and sixth orders of $\lambda$ of the normalized relative entropy compared to their second order of $\lambda$, with red, blue, green, and purple, respectively. The black dotted curve displays its numerically evaluated magnitude at an infinite order of $\lambda$. In regions where $\dot{\phi}_{0}/m_{2}r_{\rm min}$ is considerably smaller than $\sim 0.1$, each order effect is less than 100 % compared to $s^{(2)}$, indicating the validity of the perturbative calculation. For instance, the sign of the relative entropy up to the third and fifth orders of $\lambda$ changes when the red and green curves intersect the grey line, respectively. In the regions with $\dot{\phi}_{0}/m_{2}r_{\rm min}$ larger than $\sim 0.1$, the effect of each finite order is more than 100 % compared to $s^{(2)}$, and the perturbation theory is invalid. Based on the analysis of these panels, it is evident that the points where the sign flips of the relative entropy occur represent the boundaries at which the operator expansion is either valid or invalid and quantify the validity of the EFT description. This fact can be understood analytically through the order-of-magnitude evaluation for $j>2$, $|S^{(j)}/S^{(j-1)}|\sim|S^{(3)}/S^{(2)}|$. According to eq. (92), $S^{(2)}$ is non-negative, and the non-negativity breaking of $s^{(3)}$ (representing $S^{(2)}+S^{(3)}\sim 0$) indicates $|S^{(j)}/S^{(j-1)}|\sim 1$ for $j>2$, as confirmed by numerical calculation. At this juncture, it becomes unfeasible to quantitatively truncate the operator expansion beyond the third order, causing a breakdown in the operator expansion (or the EFT description).
@@
  • $\tilde{\lambda}_{4}=1$ and $\dot{\phi}_{0}/m_{2}r_{\rm min}=10^{-1}$ — Figure 3 depicts a plot similar to Fig. 2, assuming $\tilde{\lambda}_{4}=1$ and $\dot{\phi}_{0}/m_{2}r_{\rm min}=10^{-1}$ with a horizontal axis of $\tilde{\lambda}_{3}-1$. Compared to Fig. 2, the relative entropy shows almost the same qualitative behavior, except for the difference in the horizontal axis. In particular, it can be seen that, as shown in Fig. 2, the regions where the perturbation theory is valid are characterized by the point where a sign flip in the relative entropy occurs.
    Figure 3: Left panel: Normalized relative entropies, $s^{(2)}$ (black), $s^{(3)}$ (red), $s^{(4)}$ (blue), $s^{(5)}$ (green), $s^{(6)}$ (purple), and $s^{(\infty)}$ (black dotted) as a function of $\tilde{\lambda}_{3}-1$. Right panel: The size of normalized relative entropies, $s^{(3)}$ (red), $s^{(4)}$ (blue), $s^{(5)}$ (green), $s^{(6)}$ (purple), and $s^{(\infty)}$ (black dotted) compared with $s^{(2)}$ as a function of $\tilde{\lambda}_{3}-1$, respectively. The parameters are assumed to be $\tilde{\lambda}_{4}=1$ and $\dot{\phi}_{0}/m_{2}r_{\rm min}=10^{-1}$. The blue and purple curves in the right panel touch zero value. However, because of the plot’s resolution, those curves disappear in small-value of $100\times|s^{(j)}-s^{(2)}|/|s^{(2)}|$ regions.
@@ 3.2 Three scalar fields theory
We consider a scenario where a new heavy field generates the displacements of $\tilde{\lambda}_{3}$ and $\tilde{\lambda}_{4}$ from unity. Consider the following three scalar fields theory,
@@
where $\phi_{1}$ and $\phi_{2}$ are the same degrees of freedom as the scalar fields appearing in eq. (73), and $\phi_{3}$ is a real scalar field with a mass heavier than $\phi_{1}$ and $\phi_{2}$, and it behaves as a new particle in the viewpoint of the theory (73), and $V^{(3)}(\phi_{i})$ is a potential of this theory. To preserve the flatness of the inflaton potential as in eq. (74), we assume U(1) symmetry under a phase shift of $\phi_{1}+i\phi_{2}=re^{i\theta}$ in the potential $V^{(3)}(\phi_{i})$. By using the fields $r,\theta$ and $\chi:=\phi_{3}$, the action (100) can be specified at renormalizable level as follows:
@@
Upon substituting eq. (102) into eq. (46), the Euclidean effective action, up to the fourth order of the perturbative expansion, is calculated as follows:
@@
where $m^{2}_{\sigma}:=m^{2}_{2}-\beta^{2}r_{\rm min}^{4}/m_{3}^{2}$ denotes the squared mass of $\sigma$. By comparing eq. (104) with eq. (87), we find that the parameters $\tilde{\lambda}_{3}$ and $\tilde{\lambda}_{4}$ are generated from the new particle $\chi$ as follows:
@@ 3.3 An EFT of inflation generated from more general UV theories
To consider the phenomenological implications of the above inequalities (111) and (112), let us imagine the following situation. Consider the case where the Wilson coefficients $c_{2}$, $c_{3}$, etc., are determined by evaluating the observables and comparing them with experimental data. If those values of the Wilson coefficients do not satisfy the above inequalities (111) and (112), considerations based on the relative entropy suggest the following possibilities: (i) the physical phenomena are not described by the EFT (66), including the breakdown of the operator expansion; (ii) the physical phenomena are described by the EFT (66), but the loop-level contribution to the Wilson coefficients dominates; or (iii) the EFT (66) describe the phenomena, but the EFT (66) is not generated through the interaction of heavy and light fields. For each of the three cases above, we mention some possibilities. In case (i), for instance, the phenomenon is described by a theory with entirely different degrees of freedom and symmetries than the EFT (66), or the operator expansion breaks down, as seen in the sections above, meaning that UV theory, which also includes heavy degrees of freedom, is needed to describe the phenomenon. In case (ii), the UV theory of EFT (66) is the strongly coupled theory, meaning that the perturbation theory does not represent physics. In particular, this case implies that the UV theory is the strongly coupled theory, in the same sense as the $S$-matrix-based perturbative unitarity argument. Case (iii) includes, for instance, the possibility that there is no UV theory corresponding to the EFT (66), which could lead to pathological behavior, such as unitarity violation. The fact that physical models can belong to these three cases is expected to provide valuable insights into identifying models describing physical phenomena. In the next section, we will see that there are parameter regions allowed in non-Gaussianity experimental data that suggest the possibility of (i), (ii), and (iii).
@@ 4 Non-Gaussianity and non-negativity of relative entropy
Figure 4: The observationally 68% (dark blue), 95% (blue), and 99.7% (light blue) confidence regions in the $(c_{s},\tilde{c}_{3}(c_{s}^{-2}-1))$ parameter space Planck:2019kim . In the red-shaded regions corresponding to eq. (120), the non-negativity of relative entropy is violated in perturbation theory up to the third order in $X$, and the validity of the EFT (66) is lost, i.e., the operator or loop expansion becomes invalid when the EFT is generated by the interaction between heavy and light fields.
@@
where $\dot{\theta}_{0}:=\dot{\phi}_{0}/r_{\rm min}$, and these quantities are expressed in terms of parameters defined by the mass eigenstates of the heavy field in the theory with the interactions between the heavy and light fields. $m$ denotes the mass of the heavy particle (168) in that theory involving the interactions, and $\lambda_{3}$ is defined in eq. (169) of Appendix E. These parameters were introduced for convenient comparisons with previous studies Kim:2021pbr . For a detailed derivation of eq. (117), see Appendix E or Ref. Kim:2021pbr . Also, the power spectrum is expressed as,
@@
where, from the second condition of the weakly coupled theory defined in Sec. 2.2 (validity of operator expansion), we used $(c_{2}+\frac{3}{2}c_{3}\dot{\phi}_{0}^{2})^{-1}\simeq c_{2}^{-1}(1-\frac{3}{2}c_{3}\dot{\phi}_{0}^{2}/c_{2}+\cdots)$ and focused on the main contribution of the operator expansion. Figure. 4 shows allowed regions on $(c_{s},\tilde{c}_{3}(c_{s}^{-2}-1))$ plane from the current experiment Planck:2019kim . In red shaded regions in Fig. 4, the inequality (120) does not hold, and the validity of the EFT is broken. Also, by combining eqs. (116), (119) and (120), an upper bound on $f_{\rm NL}$ is obtained as follows:
@@
We see that when the EFT is generated through the weakly coupled theory (see Sec. 2.2 for its definition) with the interactions between heavy and light fields, the observationally allowed parameter region is limited. It should be emphasized that the red shaded regions in Fig. 4 are parameter regions that can be realized by the strongly coupled theory and is not physically inadmissible.
Figure 5: Contour curves of $f_{\rm NL}=\pm 1$ (black solid lines), $f_{\rm NL}=\pm 0.5$, and ±$\pm 0.1$ (black dotted lines) in the two scalar fields theory (86), shown in the plane of $m/H$ and $(\lambda_{3}-1)(\dot{\theta}_{0}/m)^{2}$, for $\dot{\theta}_{0}/H=10^{-1}$. Please also see Appendix E and Ref. Kim:2021pbr for the definition of $\lambda_{3}$ and $\lambda_{4}$. The colored dotted curves represent the perturbative unitarity bounds (124) (blue), and (125) (green). We choose $\Lambda_{\rm 5pt}=\Lambda_{\rm 6pt}=10^{5}H$ for the blue and green curves. In the red-shaded regions corresponding to eq. (122), the non-negativity of relative entropy is perturbatively violated, and the theory (86) behaves as the strongly coupled theory.
Before closing this section, we compare our method of characterizing the perturbation validity with conventional scattering amplitude-based perturbative unitarity results. Here, we compare the two methods, in particular, taking as an example the theory of two scalar fields in Sec. 3.1.2. First, from the inequality (98) up to the third order of the perturbative expansion in the relative entropy and relations presented in Appendix E, we obtain
@@
This result represents that the inequality (120), which holds for the EFT of inflation generated from general UV theories, can be confirmed using a concrete example. Next, we consider inequalities arising from the unitarity of the $S$-matrix, which has been studied in previous literature Kim:2021pbr . In theories for which the perturbative calculations are valid, inequalities arising from the unitarity of the $S$-matrix should also be satisfied at the tree-level for high-energy scatterings. As studied in Ref. Kim:2021pbr , deviations of $\lambda_{3}$ and $\lambda_{4}$ [^7] Note that $\lambda_{3}$ and $\lambda_{4}$ are the parameters defined in Appendix E for the mass eigenstate of the heavy field in the theory with the interactions between heavy and light fields and are different from $\tilde{\lambda}_{3}$ and $\tilde{\lambda}_{4}$. See Appendix E for the relation between $\tilde{\lambda}_{3,4}$ and ${\lambda}_{3,4}$. from one break the perturbative unitarity of scattering amplitudes, yielding cutoff scales of the theory if the perturbation theory represents physics faithfully. In the theory (86), the cutoff scales calculated from 5 and 6-point scatterings are provided as Kim:2021pbr
@@
From eqs. (124) and (125), we also see that the perturbative calculation of the scattering amplitude is valid for $\lambda_{3}=\lambda_{4}=1$, even in high energy regions. This fact reminds us that the non-negativity of relative entropy is unbreakable for $\lambda_{3}=\lambda_{4}=1$. It should be noted here that the scattering amplitude method characterizes the condition under which loop-level contributions do not exceed tree-level ones, whereas the relative entropy method can, in addition, characterize the validity of the operator expansion (see Sec. 2.2 for details). The $S$-matrix and the relative entropy are quantities related to the unitarity, but due to these differences, different constraints would represent the validity of the perturbation theory. From eqs. (124) and (125), we obtain
@@
As these relations depend on the cutoff scale of the theory, we show these bounds [^8] By assuming the cutoff to be $\Lambda_{\rm 5pt}\leq 10^{5}H$, and $\Lambda_{\rm 6pt}\leq 10^{5}H$, regions inside the blue or green curves are allowed from the perturbative unitarity of scattering amplitudes. in Fig. 5 for some benchmark values of the cutoff scale. Please look at the caption of Fig. 5 for more information. As shown in Fig. 5, considerations based on the relative entropy can reveal regions of the non-linear parameter where the perturbative calculations are valid and are complementary to constraints from the conventional perturbative unitarity for a given cutoff scale. It should be emphasized that while constraints from the perturbative unitarity are highly dependent on the unknown cut-off scale, the relative entropy method is not.